# WonderWorld: Interactive 3D Scene Generation from a Single Image

### v3

@@ Appendix D Further Experiment Details
where $\mathcal{U}$ is the user text input to specify the scene name. To generate the text prompt $\mathcal{T}$ for the first scene for inpainting the background layer and sky layer, we use a similar instruction to prompt the VLM (we use GPT-4V) to caption the input image.
@@
We use Prolific [^2] https://www.prolific.com/ to recruit participants for the human preference evaluation. For each experimental comparison, we recruit $204$ participants from all over the globe. We use Google forms to present the survey. The survey is fully anonymized for both the participants and the host. Participants are shown top-by-bottom bird-eye rendering images of the same layout as in Fig. 5 with randomized top-bottom orders. For the ablation study, participants are shown side-by-side images. Participants are instructed to select one from two options: “Top is more visually compelling” or “Bottom is more visually compelling” The instruction is: “Carefully compare the two images below. Which image looks better (higher quality, fewer errors) to you?”

### v4

@@ Appendix D Further Experiment Details
Real photos.
In our experiments, we use both real photos and synthetic stylized images. The following results use real photos as input: (I) “Holy Spirit Cathedral”, “Ho Chi Minh City Hall”, and “Marienplatz” in the Interactive Scene Generation section of the project website; (II) “Venice”, “Main Square”, “University Campus”, “Arc de Triomf”, “Segovia Cathedral”, “Westlake”, and “University Pathway” in the Generated Virtual World section of the project website; (III) The top example (“Venice”) and bottom right example (“Main Square”) in Fig. 1, the left example in Fig. 5, the 3rd example in Fig. 10, the 1st example in Fig. 11, the 1st example in Fig. 12, and the left example in Fig. 15.
@@
where $\mathcal{U}$ is the user text input to specify the scene name. To generate the text prompt $\mathcal{T}$ in Eq. 4 for the first scene for inpainting the background layer and sky layer, we use a similar instruction to prompt the VLM (we use GPT-4V) to caption the input image, with the difference that we also ask the VLM to generate a style prompt $\mathcal{S}$. Then, we keep using the same style prompt $\mathcal{S}$ in Eq. 4 for the whole generation process. The “scene name” above is used to prompt the LLM to generate the next scene description. The “entities” above is used as the foreground prompt $\mathcal{F}$ in Eq. 4, and the “background prompt” is used as $\mathcal{B}$ in Eq. 4.
@@
We use Prolific to recruit participants for the human preference evaluation. For each experimental comparison, we recruit $204$ participants from all over the globe. We use Google forms to present the survey. The survey is fully anonymized for both the participants and the host. Participants are shown top-by-bottom bird-eye rendering images of the same layout as in Fig. 5 with randomized top-bottom orders. For the ablation study, participants are shown side-by-side images. Participants are instructed to select one from two options: “Top is more visually compelling” or “Bottom is more visually compelling” The instruction is: “Carefully compare the two images below. Which image looks better (higher quality, fewer errors) to you?”
Figure 9: Screenshot of human study survey.
Since we compare to three baseline methods, each example forms three pairs. We use the four examples shown in Figure 5 and Figure 15, yielding $12$ pairs in total. Each participant answers all $12$ questions. We show a screenshot in Figure 9.
@@
Appendix E Additional Results
We show additional baseline comparison results in Figure 15. We show additional qualitative results in Figure 11,  10,  12. To automate generation, we also use the panoramic camera paths. We use the LLM to generate the scene names.
We show different scenes using the same input image in Figure 13, and different styles in the same virtual world in Figure 14. In Table 5, we show a time analysis of WonderWorld for generating a single extrapolated scene.
Additional ablation study on guided depth diffusion. In Table 6, we show an ablation on guided depth diffusion. Besides “w/o guided depth diffusion” which does not have any treatment to align depth estimations, we further include a heuristic-based method “Shift+Scale” which uses the least square to solve for a shift value and a scale value that transforms the estimated depth to align with the existing depth. We use the same protocol as in the baseline comparison and main paper ablation study. We report the scale-invariant root mean square error (SI-RMSE) between the estimated depth and the visible existing depth. From Table 6, we observe that our guided depth diffusion provides much better alignment than the two variants.
Figure 10: Qualitative examples. Each generated world consists of $9$ scenes. The text prompts are generated by the LLM.
Figure 11: Qualitative examples. Each generated world consists of $9$ scenes. The text prompts are generated by the LLM.
Figure 12: Qualitative examples. Each generated world consists of $9$ scenes. The text prompts are generated by the LLM.
Outpainting Layer generation Depth Normal Optim.
2.1s 2.3s 2.5s 0.8s 1.9s
Table 6: Comparison of different depth alignment methods on our examples. The metric is the scale-invariant root mean square error (SI-RMSE) between the estimated depth and the existing depth.
w/o guided depth diffusion Shift+Scale Guided depth diffusion (ours)
0.36 0.21 0.08